[figure: 9.7 cm]

RIGOROUS MAPPING OF PERIODIC ANDERSON MODEL TO KONDO LATTICE MODEL

P. Sinjukow and W. Nolting

Institute of Physics, Humboldt University Berlin, Invalidenstr.110, 10115 Berlin, Germany

We show that the Kondo lattice model can be rigorously obtained for any finite coupling constant J from the periodic Anderson model in a special limit (extended Kondo limit). The mapping goes beyond the well-known equivalence between the weak-coupling limit (J → 0) and the conventional Kondo limit of the two models. It implies a direct proof of the “large” Fermi volume for a non-magnetic Fermi-liquid state of the Kondo lattice model, which confirms in a much simpler way the result of a recent topological proof by Oshikawa [Phys. Rev. Lett. 84, 3370 (2000)]. Based on the mapping, any analytical and computational method of the periodic Anderson model, e.g. dynamical mean-field theory, can be directly applied to the Kondo lattice model.

[figure: 13.4 cm]

Subject category :

6. Theory of Magnetism

Presentation mode : poster

Corresponding author : P. Sinjukow

Address for correspondence : Institut fuer Physik Humboldt-Universitaet zu Berlin Invalidenstr.110 10115 Berlin Germany

Email address : sinjukow@physik.hu-berlin.de